# Physics Prospects with MeV Neutrino-Argon Charged Current Interactions using Enhanced Photon Detection in Future LArTPCs

### v1

@@ IV.1 Energy Reconstruction
These results come from the intrinsic anti-correlation between charge and light signals for MeV-scale $\nu_{e}$-Ar CC events, previously also observed in liquid xenon (LXe) experiments [8, 9]. The excellent energy resolution achieved by the combined calorimetry for events without knockout hadrons is further detailed in Appendix A.
@@
Energy reconstruction using the three methods is performed for all simulated samples from 5 MeV to 50 MeV under the optimistic 75 keV charge detection threshold and across all benchmark $\overline{\text{LY}}$. Since the light only distribution is non-Gaussian and both the charge-only and combined calorimetry distributions exhibit multiple secondary peaks, we use the histogram standard deviation ($\sigma_{h}$) and mean ($\bar{E}_{h}$) to characterize the energy smearing and calculate the energy resolution.
As $\nu_{e}$ energy increases,the fraction of hadron emission events rises, leading to a larger standard deviation in the reconstructed energy for all three methods and causing a degradation in energy resolution. However, above 25 MeV, the impact of undetectable binding energy from hadron emission becomes relatively small compared to the total neutrino event energy, and the energy resolution is no longer dominated by these effects. Compared to the charge-only reconstruction, the light-only reconstruction is less affected by hadron emission events, maintaining a nearly flat energy resolution as a function of true neutrino energy. This stability arises from the absence of secondary peaks, as shown in Fig. 4.
@@
For the 20 MeV $\nu_{e}$ sample shown in the bottom plot of Fig. 6, the energy resolution of the light-only method improves as $\overline{\text{LY}}$ increases. It also outperforms both charge-only and combined calorimetry reconstruction across all benchmark $\overline{\text{LY}}$, a trend that holds for $\nu_{e}$ energies above 15 MeV. The combined calorimetry method does not show significant improvement in energy resolution at higher $\overline{\text{LY}}$, as its performance remains dominated by charge calorimetry. This study demonstrates that, given sufficient light collection, light-only reconstruction can achieve superior energy resolution compared to traditional charge-based calorimetry.
@@
IV.2 Hadron Tagging
Beyond the improvements in energy reconstruction achieved through light calorimetry and dual calorimetry, further enhancement in energy resolution is expected from the ability to tag hadrons in future LArTPCs with advanced light detection systems. In particular, tagging delayed energy deposits from neutrons and their captures using timing information from the photon detection system can significantly reduce energy smearing. An example of a simulated $\nu_{e}$-Ar CC event display is shown in Fig. 7. While the variation in dE/dx is relatively small across all energy deposits, the timing domain reveals key distinguishing features. Fast timing capabilities and extensive photodetector coverage in LArTPCs can help separate energy deposits from electromagnetic activities (occurring within 10 ns) and neutron scattering activities (which extend up to a few $\mu$s). In the case of neutron capture, the capture time is primarily determined by capture on the most abundant 40Ar isotope and is expected to be on the order of $\mathcal{O}$(100) $\mu$s. The energy from the cascade gamma-rays adds up to 6.1 MeV and can be identified by the photon detection system and correlated with the primary interaction.
Figure 7: Displays of simulated energy deposits from a 50 MeV $\nu_{e}$-Ar CC event. The event vertex is located at the origin (0, 0, 0). The energy deposits are color-coded based on final state particles (top), linear energy transfer dE/dx (middle), and energy deposit time (bottom) respectively.
For this study, we assume that two common hadrons knocked out of the Ar nucleus in MeV-scale $\nu_{e}$ CC interactions—protons and $\alpha$ particles—can be tagged when their deposited charge energy exceeds the optimistic 75 keV threshold. Under this assumption, the hadron tagging efficiency is 84% for protons and 18% for $\alpha$ particles across all simulated neutrino events in the 5–50 MeV range. In the simulation, all neutrons are captured on 40Ar, and we assume all captured neutrons can be tagged in this study.
@@ IV.2 Hadron Tagging
The energy resolution after hadron tagging ($\sigma_{h}/\bar{E}_{h}$) is shown in Fig. 9 for both cases: including and excluding the tagged events. Hadron tagging significantly improves the energy resolution for the combined calorimetry reconstruction. For example, in the 25 MeV sample, the energy resolution improved from 9.4% to 5.8%. Above 25 MeV, including tagged events results in only a marginal improvement over excluding them. Meanwhile, light-only reconstruction continues to provide the best energy resolution for all $\nu_{e}$ samples above 20 MeV.
@@
Figure 8: Reconstructed ${E}_{\text{avail}}$ for 35 MeV $\nu_{e}$-Ar CC events with combined charge and light calorimetry before and after hadron tagging.
Figure 9: Reconstructed energy resolution for all events in each $\nu_{e}$ energy sample after hadron tagging. Both options of including and excluding tagged hadron events are shown. Top: reconstructed event energy resolution as a function of true ${E}_{\nu_{e}}$ from 5 to 50 MeV. Bottom: reconstructed energy resolution as a function of benchmark $\overline{\text{LY}}$ for 20 MeV $\nu_{e}$.
We note that, according to Fig. 3 and Fig. 4, one should avoid making the same correction to the reconstructed energy based on tagged hadrons in the light-only or charge-only energy reconstruction. This is due to large event-by-event fluctuations in $dE/dx$ for emitted hadrons, which affect the partitioning of energy between charge and light calorimetry according to the Birks model. Additionally, explicit charged particle identification (PID) of electrons or $\gamma$s does not further enhance the reconstructed energy resolution, as energy smearing at the deposition stage is already minimal for these particles.
IV.3 Tagging Delayed Gamma in $\nu_{e}$-Ar CC Interactions
@@
Improved vertex reconstruction: with large background rates from external and internal radioactivity, matching low-energy charge and light deposits is difficult in a large LArTPC. Vertex reconstruction is $\sim$ 1 cm for charge and limited by the light signal. Increased light collection efficiency improves vertex resolution which in turn will improve the matching efficiency. This is crucial for rejecting $\gamma$ backgrounds produced in the rock surrounding the LArTPC.
@@ V Applications to low energy physics
The energy reconstruction incorporating the light calorimetry and the assumed hadron tagging capability described in Sec IV are applied to the DSNB search as an example physics application. To refine energy smearing, the energy granularity of the Monte Carlo samples described in Sec II is enhanced by adding additional mono-energetic $\nu_{e}$ events from 5 MeV to 80 MeV in 0.5 MeV increments. For each energy step, 1000 events are generated and simulated. The final energy smearing matrices are shown in Fig. 13 in Appendix B.
@@
Fig. 10 shows the expected DSNB spectra at arbitrary exposure using the ${E}_{\text{reco},\ {Q}_{75}\ +\ {L}_{180}}$ smearing, assuming a LArTPC module with enhanced photon detection, including the hadron tagging capability as described in Sec. IV.2. Here we choose to show the event rates in arbitrary exposures to highlight the signal and background spectra and avoid explicitly deriving sensitivities for specific exposures. The vertical axis represents normalized event rates and does not correspond to actual event counts. The assumptions for the DSNB $\nu_{e}$ flux model follow those in Ref. [51]. The solar $\nu_{e}$ flux is derived from well-known nuclear cross-section measurements [54], with oscillations applied using parameters $\Delta{m}_{12}^{2}=4.86\times 10^{-5}\text{eV}^{2}$ and $\sin^{2}\theta_{12}=0.306$. Although the $\Delta{m}^{2}_{21}$ value used is slightly outdated, its impact on the results is negligible. The atmospheric $\nu_{e}$ flux is taken from Ref. [4] assuming the detector is at the SURF site.
@@
For solar backgrounds, energy smearing caused by the secondary peak from hadron knockout effects (as described in Sec IV.2) does not impact the DSNB search. The primary reconstructed energy peak from combined charge and light calorimetry, shown in Fig. 12 in Sec IV.1, provides the best opportunity to improve the search sensitivity. Conversely, for atmospheric background at the high-energy bins of the DSNB ROI, the secondary peak does affect the search sensitivity. Therefore, tagging $\nu_{e}$ CC events with knockout hadrons is crucial to mitigating smearing effects. Alternatively, one can rely on light-only energy reconstruction, which achieves better energy resolution above 20 MeV than combined calorimetry post hadron tagging, as demonstrated in Fig. 9.
To evaluate the impact of different reconstruction approaches, we apply energy smearing matrices from five reconstruction strategies: charge-only, light-only, combined calorimetry, and combined calorimetry with and without tagged hadron events, using all benchmark $\overline{\text{LY}}$.
Figure 10: Expected DSNB spectra with backgrounds at an arbitrary exposure under the best energy reconstruction strategy of combined calorimetry at a $\overline{\text{LY}}$ of 180 PE/MeV, ${E}_{\text{reco},\ {Q}_{75}\ +\ {L}_{180}}$, and exclusion of hadron tagged events. The event rate on the vertical axis uses an arbitrary unit after normalization and doesn’t represent actual number of events.
@@
Energy reconstruction using combined calorimetry provides a slight improvement in plateau sensitivity. However, when combined with the hadron tagging capability described in Sec.IV.2, it achieves the highest plateau sensitivity. Hadron tagging is particularly beneficial at higher-energy ROIs, where atmospheric background dominates. The inclusion or exclusion of tagged hadron events has minimal impact on sensitivity, so only one result is presented in Fig.11. Additionally, for $\overline{\text{LY}}$ above 180 PE/MeV, no significant improvement in sensitivity is observed for light-only, combined calorimetry, or combined calorimetry with tagged hadron events, as energy resolution is no longer the limiting factor.
@@
Further improvements in DSNB sensitivity could be achieved by tagging solar $\nu_{e}$ backgrounds using the directionality of final-state electrons in Fermi transitions, as described in Sec IV.3. This would allow statistical separation between directional solar events and the isotropic DSNB signal, enabling the ROI to be extended to lower energies and thereby increasing the DSNB event rate. Additionally, the broad energy range and baseline coverage of atmospheric neutrino oscillations in future LArTPC experiments could help constrain sub-GeV atmospheric neutrino fluxes [55].

### v2

@@ IV.1.1 Analysis A - Exclude energy deposits from neutron capture
The energy distributions obtained using three reconstruction methods in Eq. 8, Eq. 9, and Eq. 10 for 35 MeV $\nu_{e}$-Ar CC events are shown in Fig. 6. Notably, the primary peak at 34 MeV in the combined charge and light energy reconstruction (Eq. 10) is narrower than charge-only reconstruction (Eq. 8) or light-only reconstruction. This comes from the intrinsic anti-correlation between charge and light signals, which is assumed in the simulation model used in this study. This anti-correlation was previously observed in liquid xenon (LXe) experiments [16, 17]. The excellent energy resolution achieved by the combined calorimetry for events without knockout hadrons is further detailed in Appendix A.
Figure 6: Reconstructed ${E}_{\text{avail}}$ for 35 MeV $\nu_{e}$-Ar CC events using charge only (green), light only (purple), and combined charge and light energy deposits (red). The neutron capture related energy deposits are excluded during reconstruction.
Energy reconstruction using the three methods is performed for all simulated samples from 5 MeV to 50 MeV under the optimistic 75 keV charge detection threshold and across all benchmark $\overline{\text{LY}}$. Since all distributions are non-Gaussian and exhibit distinct secondary peaks, we use the histogram standard deviation ($\sigma_{h}$) and mean ($\bar{E}_{h}$) to characterize the energy smearing and calculate the energy resolution.
As true ${E}_{\nu_{e}}$ increases above 10 MeV, the fraction of hadron emission events rises (Fig. 2 bottom plot), leading to a larger standard deviation in the reconstructed energy for all three methods and causing a degradation in energy resolution. However, above a certain energy, the undetectable binding energy from hadron emission becomes relatively small compared to the total neutrino energy, and the energy resolution restores the monotonic improvement as energy increases. For the charge-only reconstruction, the turning point is around true ${E}_{\nu_{e}}$ 25 MeV. While for the light-only reconstruction, it happens slightly earlier at around true ${E}_{\nu_{e}}$ 15 MeV. Surprisingly, the light-only reconstruction also provides the best energy resolution above 15 MeV; below 15 MeV, its performance is comparable with the combined calorimetry reconstruction.
For the 20 MeV $\nu_{e}$ sample shown in the bottom plot of Fig. 7, the energy resolution of the light-only method improves as $\overline{\text{LY}}$ increases from 35 PE/MeV to 100 PE/MeV. It also outperforms both charge-only and combined calorimetry reconstruction across all benchmark $\overline{\text{LY}}$. The combined calorimetry method does not show significant improvement in energy resolution at higher $\overline{\text{LY}}$, as its performance remains dominated by charge calorimetry. This study demonstrates that, given sufficient light collection, light-only reconstruction can achieve superior energy resolution compared to traditional charge-based calorimetry.
Figure 7: Reconstructed energy resolution for all events in each $\nu_{e}$ energy sample after excluding energy deposits from neutron captures. Top: reconstructed event energy resolution as a function of true ${E}_{\nu_{e}}$ from 5 to 50 MeV. Bottom: reconstructed energy resolution as a function of benchmark $\overline{\text{LY}}$ for 20 MeV $\nu_{e}$.
IV.1.2 Analysis B - Include energy deposits from neutron capture
Beyond the improvements in energy reconstruction achieved through light calorimetry and dual calorimetry, further enhancement in energy resolution is expected from the ability to tag delayed energy deposits in future LArTPCs with advanced light detection systems. In particular, tagging delayed energy deposits from neutrons and their captures using timing information from the photon detection system can significantly reduce energy smearing. An example of a simulated $\nu_{e}$-Ar CC event display is shown in Fig. 8. While the variation in dE/dx is relatively small across all energy deposits, the timing domain reveals key distinguishing features. Fast timing capabilities and extensive photodetector coverage in LArTPCs can help separate energy deposits from electromagnetic activities (occurring within 10 ns) and neutron scattering activities (which extend up to a few $\mu$s). In the case of neutron capture, the capture time is primarily determined by capture on the most abundant 40Ar isotope and is expected to be on the order of $\mathcal{O}$(100) $\mu$s. The energy from the cascade gamma-rays adds up to 6.1 MeV and can be identified by the photon detection system and correlated with the primary interaction.
Figure 8: Displays of simulated energy deposits from a 50 MeV $\nu_{e}$-Ar CC event. The event vertex is located at the origin (0, 0, 0). The energy deposits are color-coded based on final state particles (top), linear energy transfer dE/dx (middle), and energy deposit time (bottom) respectively. In this event, the neutron has a first inelastic hadronic scattering and two gammas with energy 1.48 MeV and 0.19 MeV are produced. The neutron then goes through a series of elastic hadronic scatterings followed by a final capture on 40Ar at 1.3 $\mu$s later relative to the initial neutrino interaction. At the capture, three gammas with energy 1.92 MeV, 4.10 MeV, and 0.17 MeV are released. A 75 keV charge calorimetry detection threshold is applied in all plots.
In this second analysis, we assume neutron capture associated delayed energy deposits can be included in the physics analysis. The energy distributions obtained with this assumption with the three reconstruction methods for 35 MeV $\nu_{e}$-Ar CC events are shown in Fig. 9. The main difference to Fig. 6 is the relocation of the neutron emission events closer to the primary peak, as hinted already by Fig. 5.
@@ IV.1.2 Analysis B - Include energy deposits from neutron capture
Fig. 10 shows the updated energy resolution under the assumption neutron capture released energy deposits can be included. For $\nu_{e}$ energies below 15 MeV, the combined calorimetry method still provides the best energy resolution among the three reconstruction approaches for any benchmark $\overline{\text{LY}}$. The similar degradation in energy resolution is observed starting at 10 MeV due to the onset of hadron production, followed by a recovery at around 25 MeV. One major observed difference to the case of excluding neutron capture energy deposits in Fig. 7 is the light-only reconstruction is much less affected by hadron emission events, maintaining a nearly flat energy resolution at $\sim$4as a function of true ${E}_{\nu_{e}}$. This stability arises from the absence of secondary peaks, as shown in Fig. 4. Overall the absolute energy resolution are also better than that in Fig. 7 for each calorimetric reconstruction; the light-only reconstruction improved the most from $\sim$13to $\sim$4at true ${E}_{\nu_{e}}$ of 25 MeV.
@@
For the 20 MeV $\nu_{e}$ sample shown in the bottom plot of Fig. 10, the energy resolution of the light-only method improves as $\overline{\text{LY}}$ increases. It also outperforms both charge-only and combined calorimetry reconstruction across all benchmark $\overline{\text{LY}}$, a trend that holds for $\nu_{e}$ energies above 15 MeV. The combined calorimetry method does not show significant improvement in energy resolution at higher $\overline{\text{LY}}$, as its performance remains dominated by charge calorimetry.
@@
IV.1.3 Analysis C - Tag all hadrons and correct for the lost binding energy
The previous analysis includes all (prompt + delayed) energy deposits seen in the detector and doesn’t consider an explicit particle identification (PID) of hadrons. However, specific tagging of the hadrons using the similar photon detection system features described before can help correct for the lost binding energy associated with the hadron production. The identification of protons and $\alpha$s is possible by looking for highly localized collected charge since these hadrons have high dE/dx and are expected to only span $\sim$ 1 cm in the charge readout views in LArTPCs. An enhanced photon detection system helps to locate the region of interest in a three-dimensional sea of low-energy backgrounds so that a more efficient charge-light matching can be performed to identify these hadrons. It is promising that neutron captures can be tagged by a combined analysis of signals from both the charge and light detection systems in LArTPC through blips analysis and timing information [33, 2]. However, all these need to be further developed and demonstrated, considering potential difficulties discussed in prior literatures [47, 5]. This represents the most difficult analysis as it requires the PID of hadrons.
Building on analysis B in Sec. IV.1.2, we assume that two common hadrons knocked out of the Ar nucleus in MeV-scale $\nu_{e}$ CC interactions—protons and $\alpha$ particles—can be tagged when their deposited charge energy exceeds the optimistic 75 keV threshold. Under this assumption, the hadron tagging efficiency is 84% for protons and 18% for $\alpha$ particles across all simulated neutrino events in the 5–50 MeV range. In the simulation, all neutrons are captured on 40Ar, and we assume all captured neutrons can be tagged in this study.
@@ IV.1.3 Analysis C - Tag all hadrons and correct for the lost binding energy
The energy resolution after hadron tagging is shown in Fig. 12 for both cases: including and excluding the tagged events. Hadron tagging significantly improves the energy resolution for the combined calorimetry reconstruction. For example, in the 25 MeV sample, the energy resolution improved from 9.4% to 5.8%. Above 25 MeV, including tagged events results in only a marginal improvement over excluding them. Meanwhile, light-only reconstruction continues to provide the best energy resolution for all $\nu_{e}$ samples above 20 MeV.
@@
Figure 11: Reconstructed total available energy for 35 MeV $\nu_{e}$-Ar CC events with combined charge and light calorimetry without hadron tagging (red) same as Fig. 9, with hadron tagging: i) tagged hadron events are energy corrected and included into the primary peak (blue); ii) tagged hadron events are excluded from the sample (cyan dashed).
Figure 12: Reconstructed energy resolution for all events in each $\nu_{e}$ energy sample after hadron tagging. Both options of including (blue) and excluding (cyan dashed) tagged hadron events are shown. Top: reconstructed event energy resolution as a function of true ${E}_{\nu_{e}}$ from 5 to 50 MeV. Bottom: reconstructed energy resolution as a function of benchmark $\overline{\text{LY}}$ for 20 MeV $\nu_{e}$. In both plots, the light only calorimetry (red dotted line) and combined calorimetry (red solid line) without any hadron tagging shown in Fig 10 are overlaid for comparison.
We note that, according to Fig. 3 and Fig. 4, one should avoid making the same correction to the reconstructed energy based on tagged hadrons in the light-only or charge-only energy reconstruction. This is due to large event-by-event fluctuations in $dE/dx$ for emitted hadrons, which affect the partitioning of energy between charge and light calorimetry according to the Birks model. Additionally, explicit charged PID of electrons or $\gamma$s does not further enhance the reconstructed energy resolution, as energy smearing at the deposition stage is already minimal for these particles.
The non-uniform LY is expected to have negligible impact once precise event location is reconstructed from light signals and charge-light matching, as discussed in Appendix B.
IV.2 Tagging Delayed Gamma in $\nu_{e}$-Ar CC Interactions
@@
Improved vertex reconstruction: with large background rates from external and internal radioactivity, matching low-energy charge and light deposits is difficult in a large LArTPC. Vertex reconstruction is $\sim$ 1 cm for charge and limited by the light signal. Increased light collection efficiency improves vertex localization which in turn will improve the matching efficiency to charge signals and facilitate the reconstruction of more MeV scale activities to the cm scale. This is crucial for rejecting $\gamma$ backgrounds produced in the rock surrounding the LArTPC. The demonstration of the improved charge-light matching efficiency and the improved performance to reconstruct MeV scale energy deposition with such an enhanced light detection system will be reported in a future work.
@@ V Applications to low energy physics
The energy reconstruction incorporating the light calorimetry and the assumed hadron tagging capability described in Sec IV are applied to the DSNB search as an example physics application. To refine energy smearing, the energy granularity of the Monte Carlo samples described in Sec II is enhanced by adding additional mono-energetic $\nu_{e}$ events from 5 MeV to 80 MeV in 0.5 MeV increments. For each energy step, 1000 events are generated and simulated. The final energy smearing matrices are shown in Fig. 20 in Appendix C.
@@
Fig. 13 shows the expected DSNB spectra at 400 kt-yrs exposure using the ${E}_{\text{reco},\ {Q}_{75}\ +\ {L}_{180}}$ smearing, assuming a LArTPC module with enhanced photon detection, including the hadron tagging capability as described in Sec. IV.1.3. The assumptions for the DSNB $\nu_{e}$ flux model follow those in Ref. [52]. The atmospheric $\nu_{e}$ flux is taken from Ref. [12] assuming the detector is at the SURF site. The solar $\nu_{e}$ flux is derived from well-known nuclear cross-section measurements [55], with oscillations applied using parameters $\Delta{m}_{12}^{2}=4.86\times 10^{-5}\text{eV}^{2}$ and $\sin^{2}\theta_{12}=0.306$. Although the $\Delta{m}^{2}_{21}$ value used is slightly outdated, its impact on the results is negligible. The 3He-p endpoint 18.8 MeV does not show up in Fig. 13 and instead stops at 18 MeV because only the total reconstructed available energy is plotted and there is a 1 MeV binding energy change from the initial Ar nucleus to the K nucleus in the MARLEY event generator, as explained in Sec II. The solar neutrino backgrounds in Fig. 13 are actual shapes from the flux model after energy smearing and are part of the stacked histogram. They appear band-like because the energy resolution is very good ($\sim$1.3at 18 MeV).
@@
For solar backgrounds, energy smearing caused by the secondary peak from hadron knockout effects (as described in Sec IV.1.3) does not impact the DSNB search. The primary reconstructed energy peak from combined charge and light calorimetry, shown in Fig. 15 in Sec IV.1, provides the best opportunity to improve the search sensitivity. Conversely, for atmospheric background at the high-energy bins of the DSNB ROI, the secondary peak does affect the search sensitivity. Therefore, tagging $\nu_{e}$ CC events with knockout hadrons is crucial to mitigating smearing effects. Alternatively, one can rely on light-only energy reconstruction, which achieves better energy resolution above 20 MeV than combined calorimetry post hadron tagging, as demonstrated in Fig. 12.
To evaluate the impact of different reconstruction approaches, we apply energy smearing matrices from five reconstruction strategies: charge-only, light-only, combined calorimetry, and combined calorimetry with and without tagged hadron events, using all benchmark $\overline{\text{LY}}$. As discussed in Appendix B, the effect of non-uniform LY can be corrected with a precise knowledge of the event position, the corresponding energy smearing is expected to remain the same as the relevant light reconstruction considered above. So we don’t apply energy smearing under non-uniform light collection here.
Figure 13: Expected DSNB spectra with backgrounds at 400 kt-yrs exposure under the best energy reconstruction strategy of combined calorimetry at a $\overline{\text{LY}}$ of 180 PE/MeV, ${E}_{\text{reco},\ {Q}_{75}\ +\ {L}_{180}}$, and exclusion of hadron tagged events.
@@
Energy reconstruction using combined calorimetry provides a slight improvement in plateau sensitivity. However, when combined with the hadron tagging capability described in Sec.IV.1.3, it achieves the highest plateau sensitivity. Hadron tagging is particularly beneficial at higher-energy ROIs, where atmospheric background dominates. The inclusion or exclusion of tagged hadron events has minimal impact on sensitivity, so only one result is presented in Fig.14. Additionally, for $\overline{\text{LY}}$ above 180 PE/MeV, no significant improvement in sensitivity is observed for light-only, combined calorimetry, or combined calorimetry with tagged hadron events, as energy resolution is no longer the limiting factor.
@@
Further improvements in DSNB sensitivity could be achieved by tagging solar $\nu_{e}$ backgrounds using the directionality of final-state electrons in Fermi transitions, as described in Sec IV.2. This would allow statistical separation between directional solar events and the isotropic DSNB signal, enabling the ROI to be extended to lower energies and thereby increasing the DSNB event rate. Additionally, the broad energy range and baseline coverage of atmospheric neutrino oscillations in future LArTPC experiments could help constrain sub-GeV atmospheric neutrino fluxes [56].